OpenStudy (anonymous):

For the system of linear equations below, use graphing to determine in which quadrant the solution lies. x+2y=-2 -2x+y=14

OpenStudy (anonymous):

Well you have to put the two equations in y=mx + b form.

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

x+2y=-2 -x -x 2y = -x - 2 /2 /2 /2 y = (-x/2) - 1

OpenStudy (anonymous):

So would that be in quadrant IV?

OpenStudy (anonymous):

Let me see

OpenStudy (anonymous):

No the solution would be in quadrant II

OpenStudy (anonymous):

Okay thanks :-)

OpenStudy (anonymous):

No prob :)